### v2

# The universal family of punctured Riemann surfaces is Stein

Abstract.
We show that for any pair of integers $n\geq 1$ and $g\geq 0$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold.
Key words and phrases:
Riemann surface, Teichmüller space, universal family, Stein manifold, Oka manifold
2020 Mathematics Subject Classification:
Primary 32G15; secondary 32Q28, 32Q56
1. Introduction
[…]
(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$.
In this paper we prove the following result.
Theorem 1.1.
If $n\geq 1$ then the universal family $V(g,n)$ is a Stein manifold.
The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [22, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [21]). Note also that $V(g,0)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,0)$.
The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1.1), corresponding to the punctures of $V(g,n)$, are called canonical sections in the literrature. The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [19] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$. See also [20]. Earle and Kra proved in [7, p. 50] that for any $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
Theorem 1.1 has interesting consequences for function theory and complex geometry. By classical results of Remmert, Bishop, and Narasiman (see [9, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$ (which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$). When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ by the optimal embedding theorems of Eliashberg and Gromov [8] and Schürmann [24]. Since every fibre of $V(g,n)$ is a noncompact Riemann surface of finite conformal type, and hence is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
Problem 1.2.
Does $V(g,n)$ for $n\geq 1$ admit a proper holomorphic embedding in some $\mathbb{C}^{N}$ which is algebraic on every fibre?
Except in the trivial cases, $T(g,n)$ is not affine algebraic, and it is not reasonable to expect that $V(g,n)$ is affine algebraic.
Since $T(g,n)$ is contractible, the holomorphic submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$. The following is an immediate corollary to this observation, Theorem 1.1, and the main result of Oka theory [9, Theorem 5.4.4]. For Oka manifolds, see [9, Chap. 5] and the surveys [10, 12, 13].
Corollary 1.3.
Let $\pi:V(g,n)\to T(g,n)$ be as above, $n\geq 1$, and let $Y$ be an Oka manifold. There is a holomorphic map $V(g,n)\to Y$ in every homotopy class. Furthermore, a holomorphic map $M_{t}\to Y$, $t\in T(g,n)$, from any fibre of $V(g,n)$ extends to a holomorphic map $V(g,n)\to Y$. More generally, given a closed complex subvariety $T^{\prime}$ of $T(g,n)$, every continuous map $F_{0}:V(g,n)\to Y$ which is holomorphic on $\pi^{-1}(T^{\prime})$ is homotopic to a holomorphic map $F:V(g,n)\to Y$ by a homotopy which is fixed on $\pi^{-1}(T^{\prime})$.
If $n\geq 1$ and $(g,n)\neq(0,1)$ then $V(g,n)$ is not simply connected, and its homotopy type is that of a finite bouquet of circles. In this case, homotopically nontrivial maps $V(g,n)\to Y$ exist whenever the manifold $Y$ is not simply connected. This gives the following corollary. The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$.
Corollary 1.4.
If $n\geq 1$, $(g,n)\neq(0,1)$, and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre. In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [9, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
In this connection, we mention that in [11] the Oka principle was established for families of maps from very general families of open Riemann surfaces $(M,J_{t})$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space, which is not a complex manifold in general. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [11, Theorem 1.6] are merely continuous or smooth in the parameter. In [14] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds. Here, a continuous family $\{J_{t}\}_{t\in T}$ of Stein structures on a smooth manifold $X$ said to be tame if the holomorphically convex hulls of any compact subset $K\subset X$ are upper semicontinuous with respect to the parameter $t\in T$. This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
2. Proof of Theorem 1.1
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ in (1.1), Theorem 1.1 is an immediate consequence of the following more general result with arbitrary Stein manifold as the base.
Theorem 2.1.
Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
The conclusion fails if the fibres of $\pi$ have complex dimension >1>1 or the sections $s_{i}$ are not holomorphic. In such case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$. Note that Stein complements of smooth complex hypersurfaces in compact Kähler manifolds have recently been studied by Höring and Peternell [18] where the reader can find references to earlier works. In our case, $Z$ is not compact unless $X$ is a point.
Proof.
Let $\pi:Z\to X$ be as in the theorem. We may assume that $X$ is connected. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
Recall the following result (see Grauert and Remmert [16, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [16, Chap. A.3.5]). Hence, is suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a single holomorphic section $s:X\to Z$ is a Stein manifold.
By a theorem of Siu [25], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [16, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [15] will then imply that $Z\setminus H$ is Stein.
Fix a point $x_{0}\in X$ and set $z_{0}=s(x_{0})\in H\subset Z$. Since $\pi:Z\to X$ is a holomorphic submersion with compact one-dimensional fibres, it is a smooth fibre bundle whose fibre $M$ is a compact smooth surface. In particular, there is a neighbourhood $X_{0}\subset X$ of $x_{0}$ such that the restricted bundle $Z|X_{0}=\pi^{-1}(X_{0})\to X_{0}$ can be smoothly identified with the trivial bundle $X_{0}\times M\to X_{0}$. In this identification, $z_{0}=(x_{0},p_{0})$ with $p_{0}\in M$. Since $\phi$ tends to $\infty$ along $H$, there are small smoothly bounded open discs $D\Subset D^{\prime}\Subset M$ with $p_{0}\in D$ such that
(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
$u_{0}<\phi(x_{0},\cdotp)$ on $bO=bD$ and $u_{0}>\phi(x_{0},\cdotp)$ on $bD^{\prime}$.
Shrinking the neighbourhood $X_{0}\subset X$ of $x_{0}$ if necessary, the following conditions hold for every $x\in X_{0}$, where we use the smooth fibre bundle isomorphism $Z|X_{0}\cong X_{0}\times M$:
  1. (a)
    $s(x)\in D$,
  2. (b)
    the function $u(x,\cdotp)=u_{0}$ is strongly subharmonic on $O$ in the complex structure on $Z_{x}\cong M$,
  3. (c)
    $u(x,\cdotp)<\phi(x,\cdotp)$ on $bO=bD$, and
  4. (d)
    $u(x,\cdotp)>\phi(x,\cdotp)$ on $bD^{\prime}$.
Condition (b) holds since being strongly subharmonic on a compact subset is a stable property under small smooth deformations of the complex structure. We define a function $\rho_{0}:(X_{0}\times M)\setminus H\to\mathbb{R}_{+}$ by taking for every $x\in X_{0}$:
$\rho_{0}(x,p)=\begin{cases}\phi(x,p),&p\in D\setminus\{s(x)\};\\ \max\{\phi(x,p),u(x,p)\},&p\in D^{\prime}\setminus D;\\ u(x,p),&p\in M\setminus D^{\prime}.\end{cases}$
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [9, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H=\pi^{-1}(X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$. Let $\{\chi_{j}\}_{j}$ be a smooth partition of unity on $X$ with compact supports $\mathrm{supp}(\chi_{j})\subset X_{j}$ for each $j$. Set
$\rho=\sum_{j=1}^{\infty}\chi_{j}\rho_{j}:Z\setminus H\to\mathbb{R}_{+}.$
By the construction, the restriction of $\rho$ to each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X)$ is strongly subharmonic, and there is an open neighbourhood $U_{0}\subset U\subset Z$ of $H$ such that $\rho=\phi$ holds on $U_{0}\setminus H$. In particular, $\rho$ is strongly plurisubharmonic on $U_{0}\setminus H$.
Note that for every compact set $K\subset X$, the set $\pi^{-1}(K)\setminus U_{0}\subset Z\setminus H$ is compact. Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that $\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$ is a strongly plurisubharmonic exhaustion function. Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by $(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)$ for $z\in Z$ and $\xi\in T_{z}Z$, $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if $(dd^{c}\rho)(z,\xi\wedge J\xi)>0$ for every tangent vector $0\neq\xi\in T_{z}Z$. (Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since the function $\rho$ constructed above is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that
$(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{if $z\in Z\setminus H$ and $0\neq\xi\in\ker d\pi_{z}$.}$
Hence, the eigenvectors of the Levi form $(dd^{c}\rho)(z,\cdotp)$ associated to non-positive eigenvalues lie in a closed cone $C_{z}\subset T_{z}Z$ which intersects $\ker d\pi_{z}$ only in the origin. It follows that if $\tau:X\to\mathbb{R}$ is such that $dd^{c}\tau>0$ is sufficiently big on $T_{x}X$, $x=\pi(z)$, then $dd^{c}\rho+dd^{c}(\tau\circ\pi)>0$ on $T_{z}Z$. Furthermore, the estimates are uniform on the compact set $\pi^{-1}(K)\setminus U_{0}$, where $U_{0}\subset Z$ is a neighbourhood of $H$ as above such that $dd^{c}\rho>0$ on $U_{0}\setminus H$. To see that $\tau$ can be chosen such that $dd^{c}\tau$ grows as fast as desired, note that if $h:\mathbb{R}\to\mathbb{R}$ is a $\mathscr{C}^{2}$ function then for each point $x\in X$ and tangent vector $\xi\in T_{x}X$ we have that
$dd^{c}(h\circ\tau)(x,\xi\wedge J\xi)=h^{\prime}(\tau(x))\,(dd^{c}\tau)(x,\xi\wedge J\xi)+h^{\prime\prime}(\tau(x))\,\big(|d\tau(x,\xi)|^{2}+|d\tau(x,J\xi|^{2}\big).$
Hence, if $\tau$ is a strongly plurisubharmonic exhaustion function on $X$ and the function $h:\mathbb{R}\to\mathbb{R}$ is chosen such that $h^{\prime\prime}\geq 0$ and $h^{\prime}$ grows sufficiently fast, then $dd^{c}(h\circ u)$ also grows as fast as desired.
This completes the proof of Theorem 2.1, and it also proves Theorem 1.1. ∎
A minor modification of the proof of Theorem 2.1 gives the following more general result.
Theorem 2.2.
Assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$. Then, the domain $\Omega=Z\setminus H$ is Stein.
Sketch of proof.
The restricted projection $\pi:H\to X$ is a proper holomorphic map. Since $H$ is of pure codimension one and does not contain any fibre of $\pi$, $\pi:H\to X$ is a finite holomorphic map (a branched holomorphic cover). The proper mapping theorem of Remmert [23] (see also Chirka [6, p. 29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [16, Theorem 1 (d), p. 125]. By Siu’s theorem [25], $H$ admits a basis of open Stein neighbourhoods $U\subset Z$. By the argument in the proof of Theorem 2.1, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [16, Theorem 5, p. 129]. The proof can now be completed by a similar argument as in the proof of Theorem 2.1, and we leave further details to the reader. ∎
Acknowledgements. Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia. I wish to thank Finnur Lárusson for posing the question answered by Theorem 1.1 in a private communication.