### v4

# The universal family of punctured Riemann surfaces is Stein

Abstract.
We show that the universal Teichmüller family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold for any $n>0$. We describe its basic function theoretic properties and pose several challenging questions. We show in particular that the space of fibrewise algebraic functions on the universal family is dense in the space of holomorphic functions, and there is a fibrewise algebraic map of the universal family in a Euclidean space which is an embedding over any given relatively compact domain in the Teichmüller space. We also obtain a relative Oka principle for holomorphic fibrewise algebraic maps of the universal family to any flexible algebraic manifold.
Key words and phrases:
Riemann surface, Teichmüller space, universal family, Stein manifold, Oka manifold
2020 Mathematics Subject Classification:
Primary 32G15; secondary 32Q28, 32Q56
1. Introduction
[…]
(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$. If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [39, Theorem 5.4.3]).
In this paper, we establish several function theoretic properties of the universal family $V(g,n)$. Our first main result is the following.
Theorem 1.1.
The Teichmüller family $V(g,n)$ for $n\geq 1$ is a Stein manifold.
This is a special case of Theorems 2.1 and 2.2 given in the sequel. The result is obvious if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ is biholomorphic to $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [39, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ since the Kobayashi metric and the Carathéodory metric on it differ (see Marković [38]), while they agree on a convex domain (see Lempert [37]). Note also that $\widehat{V}(g,n)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,n)$.
We provide two proofs of Theorem 1.1. The first more elementary one in Sect. 2 shows that $V(g,n)$ admits a strongly plurisubharmonic exhaustion function, hence is Stein by a theorem of Grauert [26]. The second one follows from Theorem 3.1 which shows that holomorphic fibrewise algebraic functions on $V(g,n)$ separate points and provide holomorphic convexity. The proof uses Grauert’s theorem on coherence of direct images of coherent analytic sheaves and other techniques of complex geometry.
Theorem 1.1 has several interesting consequences. By classical results of Remmert, Bishop, and Narasiman (see [19, Theorem 2.4.1] and the references therein), $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$. When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [16] and Schürmann [42], or the exposition in [19, Secs. 9.3–9.4]).
Since the Teichmüller space $T(g,n)$ is contractible, the submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class. The following is a corollary to this observation, Theorem 1.1, and the main result of Oka theory [19, Theorem 5.4.4]. (See also the surveys [20, 21, 23].)
Corollary 1.2.
Let $\pi:V(g,n)\to T(g,n)$ be as above, $n\geq 1$, and let $Y$ be an Oka manifold. There is a holomorphic map $V(g,n)\to Y$ in every homotopy class. Furthermore, a holomorphic map $M_{t}\to Y$, $t\in T(g,n)$, from any fibre of $V(g,n)$ extends to a holomorphic map $V(g,n)\to Y$. More generally, given a closed complex subvariety $T^{\prime}$ of $T(g,n)$, every continuous map $F_{0}:V(g,n)\to Y$ which is holomorphic on $\pi^{-1}(T^{\prime})$ is homotopic to a holomorphic map $F:V(g,n)\to Y$ by a homotopy which is fixed on $\pi^{-1}(T^{\prime})$.
If $n\geq 1$ and $(g,n)\neq(0,1)$ then $V(g,n)$ is not simply connected, and its homotopy type is that of a finite bouquet of circles. In this case, homotopically nontrivial maps $V(g,n)\to Y$ exist whenever the manifold $Y$ is not simply connected. This gives the following corollary. The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$. The result obviously holds for g=0,n=1g=0,\ n=1 since $V(0,1)=\mathbb{C}$.
Corollary 1.3.
If $n\geq 1$ and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre. In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [19, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
Since every fibre of $V(g,n)$ is an affine algebraic curve, it is natural to expect that $V(g,n)$ admits holomorphic functions which are algebraic on every fibre. This is indeed the case. We show in Theorem 3.4 that the algebra $\mathscr{A}(V(g,n))$ of fibrewise algebraic functions on $V(g,n)$ is dense in the algebra $\mathscr{O}(V(g,n))$ of holomorphic functions in the compact-open topology. Furthermore, functions in $\mathscr{A}(V(g,n))$ provide a proper embedding of $V(g,n)$ in a Euclidean space over any relatively compact domain in $T(g,n)$ (see Theorem 3.1), and fibrewise algebraic maps of $V(g,n)$ to any flexible algebraic manifold satisfy a relative Oka principle (see Theorem 4.1). The following problem remains open.
Problem 1.4.
Does $V(g,n)$ for $n\geq 1$ admit a proper holomorphic embedding in some $\mathbb{C}^{N}$ which is algebraic on every fibre?
The fact that $V(g,n)$ for $n\geq 1$ is homotopy equivalent to a bouquet of circles implies that every complex vector bundle on $V(g,n)$ is topologically trivial. Since $V(g,n)$ is Stein, the Oka–Grauert principle (see Grauert [25] or [19, Theorem 3.2.1]) implies the following.
Proposition 1.5.
Every holomorphic vector bundle on $V(g,n)$ for $n\geq 1$ is holomorphically trivial.
This fails in the algebraic category, even on a single fibre. In fact, the tangent bundle of a finitely punctured compact Riemann surface is not algebraically trivial in general.
Corollary 1.6.
Assume that $g\geq 0$ and $n\geq 1$.
  1. (a)
    There exists a nowhere vanishing holomorphic vector field $\xi$ on $V(g,n)$ which is tangent to the fibres of the projection $\pi:V(g,n)\to T(g,n)$, that is, $d\pi(\xi)=0$.
  2. (b)
    With $\xi$ as in (a), there exists a holomorphic $1$-form $\theta$ on $V(g,n)$ satisfying $\langle\theta,\xi\rangle=1$. In particular, $\theta$ is nowhere vanishing on the tangent bundle to any fibre of $\pi$.
Proof.
Part (a) follows by applying Proposition 1.5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$. To see (b), consider the following short exact sequence of vector bundles over $V(g,n)$:
$0\longrightarrow\ker d\pi\lhook\joinrel\relbar\joinrel\rightarrow TV(g,n)\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}H:=TV(g,n)/\ker\pi\longrightarrow 0.$
Here, $TV(g,n)$ denote the tangent bundle of $V(g,n)$. By Cartan’s Theorem B the sequence splits, i.e. there is a holomorphic vector bundle injection $\sigma:H\hookrightarrow TV(g,n)$ such that $\alpha\circ\sigma=\mathrm{Id}_{H}$. Hence, $TV(g,n)=\ker d\pi\oplus\sigma(H)=\mathbb{C}\xi\oplus\sigma(H)$, where $\xi$ is as in part (a). The unique holomorphic $1$-form $\theta$ on $V(g,n)$ satisfying $\langle\theta,\xi\rangle=1$ and $\xi=0$ on $\sigma(H)$ clearly satisfies part (b). ∎
By the Gunning–Narasimhan theorem [31], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 1.6 (b), the following is a natural question.
Problem 1.7.
Let $n\geq 1$. Is there a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to every fibre of $V(g,n)$ is an immersion?
By [17, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points. The problem is to find $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$. Note that [22, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre is a holomorphic immersion. Problem 1.7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 1.6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$. However, this is not the only problem. Since the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$ (see Hubbard [34, 35] and Earle and Kra [15, p. 50] for a precise description of sections of $\widehat{V}(g,n)$ and $V(g,n)$), there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
Another interesting question is whether the Riemann surfaces $M_{t}$ in the Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$. For background, see [40] and [4]. By [22, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter. It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature. Each single surface in the family can be made such by [7, 6].
It is natural to wonder whether the results of this paper extend to infinite dimensional Teichmüller families. The analogous problem with continuous or smooth dependence of the complex structures and the holomorphic maps on the parameter was studied in [22] for very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t$ in a topological space $T$. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. Under mild assumptions on the family $\{J_{t}\}_{t\in T}$, $J_{t}$-holomorphic functions on $(M,J_{t})$ satisfy the Runge approximation theorem with continuous or smooth dependence on $t\in T$ [22, Theorem 1.1]. For a class of parameter spaces including finite CW complexes we also have the Oka principle for continuous or smooth families of holomorphic maps from $(M,J_{t})$ to any Oka manifold [22, Theorem 1.6]. These results were extended in [24] to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds.
2. Proof of Theorem 1.1
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ (1.1), Theorem 1.1 is an immediate consequence of the following result with arbitrary Stein manifold as the base.
Theorem 2.1.
Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected fibres, and $s_{1},\ldots,s_{n}:X\to Z$ $(n\geq 1)$ are holomorphic sections with pairwise disjoint images. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
The assumption that the sections $s_{1},\ldots,s_{n}$ have pairwise disjoint images is inessential and is imposed only for convenience of the proof; see Theorem 2.2 for a more general result. The conclusion fails if the fibres of $\pi$ have complex dimension >1>1, or if the sections $s_{i}$ are not holomorphic. In such a case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$. Note that Stein complements of smooth complex hypersurfaces in compact Kähler manifolds have recently been studied by Höring and Peternell [32] where the reader can find references to earlier works. In our case, $Z$ is not compact unless $X$ is a point.
Proof of Theorem 2.1.
Let $\pi:Z\to X$ be as in the theorem. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact connected Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
Recall the following result (see Grauert and Remmert [27, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [27, Chap. A.3.5]). Hence, it suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a holomorphic section $s:X\to Z$ is a Stein manifold.
By a theorem of Siu [44], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold [27, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [26] will then imply that $Z\setminus H$ is Stein.
Fix a point $x_{0}\in X$ and set $z_{0}=s(x_{0})\in H\subset Z$. Since $\pi:Z\to X$ is a holomorphic submersion with compact one-dimensional fibres, it is a smooth fibre bundle whose fibre $M$ is a compact smooth surface. In particular, there is a neighbourhood $X_{0}\subset X$ of $x_{0}$ such that the restricted bundle $Z|X_{0}=\pi^{-1}(X_{0})\to X_{0}$ can be smoothly identified with the trivial bundle $X_{0}\times M\to X_{0}$. In this identification, $z_{0}=(x_{0},p_{0})$ with $p_{0}\in M$. Since $\phi$ tends to $\infty$ along $H$, there are small smoothly bounded open discs $D\Subset D^{\prime}\Subset M$ with $p_{0}\in D$ such that
(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
$u_{0}<\phi(x_{0},\cdotp)$ on $bO=bD$ and $u_{0}>\phi(x_{0},\cdotp)$ on $bD^{\prime}$.
Shrinking the neighbourhood $X_{0}\subset X$ of $x_{0}$ if necessary, the following conditions hold for every $x\in X_{0}$, where we use the smooth fibre bundle isomorphism $Z|X_{0}\cong X_{0}\times M$:
  1. (a)
    $s(x)\in D$,
  2. (b)
    the function $u(x,\cdotp)=u_{0}$ is strongly subharmonic on $O$ in the complex structure on $Z_{x}\cong M$,
  3. (c)
    $u(x,\cdotp)<\phi(x,\cdotp)$ on $bO=bD$, and
  4. (d)
    $u(x,\cdotp)>\phi(x,\cdotp)$ on $bD^{\prime}$.
Condition (b) holds since being strongly subharmonic on a compact subset is a stable property under small smooth deformations of the complex structure. We define a function $\rho_{0}:(X_{0}\times M)\setminus H\to\mathbb{R}_{+}$ by taking for every $x\in X_{0}$:
$\rho_{0}(x,p)=\begin{cases}\phi(x,p),&p\in D\setminus\{s(x)\};\\ \max\{\phi(x,p),u(x,p)\},&p\in D^{\prime}\setminus D;\\ u(x,p),&p\in M\setminus D^{\prime}.\end{cases}$
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [19, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$. Let $\{\chi_{j}\}_{j}$ be a smooth partition of unity on $X$ with compact supports $\mathrm{supp}(\chi_{j})\subset X_{j}$ for each $j$. Set
$\rho=\sum_{j=1}^{\infty}\chi_{j}\rho_{j}:Z\setminus H\to\mathbb{R}_{+}.$
By the construction, the restriction of $\rho$ to each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X)$ is strongly subharmonic, and there is an open neighbourhood $U_{0}\subset U\subset Z$ of $H$ such that $\rho=\phi$ holds on $U_{0}\setminus H$. In particular, $\rho$ is strongly plurisubharmonic on $U_{0}\setminus H$. Note that for every compact set $K\subset X$, the set $\pi^{-1}(K)\setminus U_{0}\subset Z\setminus H$ is compact. Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that
$\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$
is a strongly plurisubharmonic exhaustion function, thus proving the theorem. Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by
$(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)\ \ \text{for $z\in Z$ and $\xi\in T_{z}Z$},$
the function $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if
$(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{for every tangent vector $0\neq\xi\in T_{z}Z$}.$
(Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since $\rho$ is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that
$(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{if $z\in Z\setminus H$ and $0\neq\xi\in\ker d\pi_{z}$.}$
Hence, the eigenvectors of $(dd^{c}\rho)(z,\cdotp)$ associated to non-positive eigenvalues lie in a closed cone $C_{z}\subset T_{z}Z$ which intersects $\ker d\pi_{z}$ only in the origin. It follows that if $\tau:X\to\mathbb{R}$ is such that $dd^{c}\tau>0$ is sufficiently big on $T_{x}X$ where $x=\pi(z)$, then $dd^{c}\rho+dd^{c}(\tau\circ\pi)>0$ on $T_{z}Z$. Furthermore, the estimates are uniform on the compact set $\pi^{-1}(K)\setminus U_{0}$, where $U_{0}\subset Z$ is a neighbourhood of $H$ such that $dd^{c}\rho>0$ on $U_{0}\setminus H$. To see that $\tau$ can be chosen such that $dd^{c}\tau$ grows as fast as desired, note that if $h:\mathbb{R}\to\mathbb{R}$ is a $\mathscr{C}^{2}$ function then for each point $x\in X$ and vector $\xi\in T_{x}X$ we have that
$dd^{c}(h\circ\tau)(x,\xi\wedge J\xi)=h^{\prime}(\tau(x))\,(dd^{c}\tau)(x,\xi\wedge J\xi)+h^{\prime\prime}(\tau(x))\,\big(|d\tau(x,\xi)|^{2}+|d\tau(x,J\xi)|^{2}\big).$
Hence, if $\tau$ is a strongly plurisubharmonic exhaustion function on $X$ and the function $h:\mathbb{R}\to\mathbb{R}$ is chosen such that $h^{\prime\prime}\geq 0$ and $h^{\prime}$ grows sufficiently fast, then $dd^{c}(h\circ\tau)$ also grows as fast as desired. ∎
A minor modification of the proof of Theorem 2.1 gives the following more general result.
Theorem 2.2.
Assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected one-dimensional fibres, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$. Then, the domain $\Omega=Z\setminus H$ is Stein.
Proof.
The hypotheses imply that $\pi|_{H}:H\to X$ is a finite holomorphic map (a branched holomorphic cover). The proper mapping theorem of Remmert [41] (see also [12, p. 29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $\dim X$, hence $\pi(H)=X$ since $X$ is connected. By [27, Theorem 1 (d), p. 125], the subvariety $H$ is Stein. A theorem of Siu [44] implies that $H$ has a basis of open Stein neighbourhoods $U\subset Z$. By the argument in the proof of Theorem 2.1, the ideal sheaf of $H$ is a principal ideal sheaf, so the manifold $U\setminus H$ is Stein for any Stein neighbourhood $U\subset Z$ of $H$ [27, Theorem 5, p. 129]. The proof can now be completed by a similar argument as in the proof of Theorem 2.1, and we leave the details to the reader. ∎
Remark 2.3.
Theorem 2.2 still holds if the fibres of the projection $\pi:Z\to X$ are not necessarily connected provided that the subvariety $H$ intersects every connected component of each fibre $Z_{x}=\pi^{-1}(x),\ x\in X$, but it does not contain any such component. The result also holds if $X$ is a Stein space, $Z$ is a complex space, $H\subset Z$ is a complex hypersurface whose ideal sheaf is principal, and the other conditions on the submersion $\pi:Z\to X$ and $H$ are satisfied.
3. Fibrewise algebraic functions on $V(g,n)$
In this section, we show that the universal family $V(g,n)$ of $n$-punctured compact Riemann surfaces of genus $g$ for $n\geq 1$ has a large algebra of holomorphic functions which are algebraic on every fibre of the Teichmüller projection $\pi:V(g,n)\to T(g,n)$. Indeed, such functions are dense in the algebra of holomorphic functions in the compact-open topology; see Theorem 3.4.
We shall consider the more general situation given in Theorem 2.2. Thus, we assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected one-dimensional fibres, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$. The domain $\Omega=Z\setminus H$ is Stein by Theorem 2.2, and every fibre $\Omega_{x}$ $(x\in X)$ is a finitely punctured compact Riemann surface, hence affine. Let $\mathscr{A}(\Omega)$ denote the $\mathbb{C}$-algebra of holomorphic functions on $\Omega$ which are algebraic on every fibre $\Omega_{x}$, $x\in X$. Clearly, $\mathscr{A}(\Omega)$ is a subalgebra of the algebra $\mathscr{O}(\Omega)$ of holomorphic functions.
Theorem 3.1.
(Assumptions as above.) Given a relatively compact domain $U\Subset X$, there exist finitely many functions $f_{1},\ldots,f_{n}\in\mathscr{A}(\Omega)$ such that the map $F:\Omega\to X\times\mathbb{C}^{n}$ given by
(3.1) $F(z)=(\pi(z),f_{1}(z),\ldots,f_{n}(z)),\quad z\in\Omega$
induces a proper embedding $\Omega_{U}:=\Omega\cap\pi^{-1}(U)\to U\times\mathbb{C}^{n}$.
A map $F$ satisfying the conclusion of theorem is called a proper embedding over $U$. The same result holds if $X$ is a Stein space, $Z$ is a complex space, and the other conditions on the submersion $\pi:Z\to X$ and the hypersurface $H\subset Z$ hold; in particular, the ideal sheaf of $H$ is principal. Theorem 3.1 clearly implies that holomorphic fibrewise algebraic functions on $\Omega$ separate points and provide holomorphic convexity, thus providing another proof that $\Omega$ (and in particular the universal family $V(g,n)$) is a Stein manifold.
Proof.
Let $L\to Z$ denote the holomorphic line bundle determined by the divisor $H\subset X$, and let $\sigma_{0}:Z\to L$ be a holomorphic section whose zero divisor equals $H$. (The existence of such a section is tautological from the construction of $L$.) Since the fibre $Z_{x}=\pi^{-1}(x)$ is a projective curve for every $x\in X$, Serre’s GAGA principle [43] implies that the restricted line bundle $L_{x}:=L|Z_{x}\to Z_{x}$ is algebraic, and every holomorphic section of $L_{x}$ over $Z_{x}$ is algebraic. It follows that for any holomorphic section $\sigma:Z\to L$, the quotient $f=\sigma/\sigma_{0}$ is a well defined function in $\mathscr{A}(\Omega)$. The restriction of $f$ to $\Omega_{x}$ is algebraic and has an effective pole at every end of $\Omega_{x}$ (that is, a point of $H_{x}=H\cap Z_{x}$) at which $\sigma$ does not vanish. Note that the line bundle $L_{x}\to Z_{x}$ is associated to the effective divisor supported on $H_{x}$ in which the multiplicity of a point $z\in H_{x}$ is the intersection number of $H$ with $Z_{x}$ at $z$. In particular, $L_{x}$ is an ample line bundle on $Z_{x}$ for every $x\in X$. Hence, some tensor power $L_{x}^{\otimes d}\to Z_{x}$ for $d=d(x)>0$ is very ample, so it admits finitely many holomorphic sections $\sigma_{1,x},\cdots,\sigma_{n,x}:Z_{x}\to L_{x}$ such that the map
(3.2) $Z_{x}\ni z\longmapsto\big[\sigma_{0,x}^{d}(z):\sigma_{1,x}(z):\cdots:\sigma_{n,x}(z)\big]\in\mathbb{CP}^{n}$
is a holomorphic embedding. Here, $\sigma_{0}^{d}$ denotes the $d$-th power of $\sigma_{0}$, a section of the line bundle $L^{\otimes d}\to Z$. Note that $\sigma_{i,x}/\sigma_{0,x}^{d}$ is a regular algebraic function on $\Omega_{x}$ for every $i=1,\ldots,n$.
To prove the theorem, we will show that for $d=d(x)>0$ big enough, every holomorphic section $\sigma_{x}:Z_{x}\to L_{x}^{\otimes d}$ extends to a holomorphic section $\sigma:Z\to L^{\otimes d}$. Assume for a moment that this holds true. Extending all section $\sigma_{i,x}$ in (3.2) for $i=1,\ldots,n$ to $Z$, the map in (3.2) is an embedding for any base point in an open neighbourhood $U_{1}\subset X$ of $x$. Hence, the map $F$ of the form (3.1) with the components $f_{i}=\sigma_{i}/\sigma_{0}^{d}\in\mathscr{A}(\Omega)$ is a proper embedding over $U_{1}$. Indeed, for any pair of points $x\in U_{1}$ and $z\in H_{x}$, at least one of the sections $\sigma_{i,x}$ $(i=1,\ldots,n)$ does not vanish at $z$. Since $\sigma_{0}(z)=0$, it follows that $f_{i}=\sigma_{i}/\sigma_{0}^{d}$ has an effective pole at $z$, which implies properness. Repeating the same construction at other points $x\in\overline{U}$ (for different values of $d=d(x)\in\mathbb{N}$) and assembling the resulting functions as components of a map (3.1) yields a proper embedding over $U$ as in the theorem.
It remains to explain why every holomorphic section $\sigma_{x}:Z_{x}\to L_{x}^{\otimes d}$ for $d>0$ big enough extends to a holomorphic section $\sigma:Z\to L^{\otimes d}$. Let $\mathscr{L}\to Z$ denote the sheaf of holomorphic sections of the line bundle $L\to Z$; note that $\mathscr{L}$ is a coherent $\mathscr{O}_{Z}$-analytic sheaf. Since the holomorphic projection $\pi:Z\to X$ is surjective and proper, the direct image sheaf $\pi_{*}\mathscr{L}$ is a coherent analytic sheaf on $X$ by Grauert’s direct image theorem [28, p. 207]. A section of $\pi_{*}\mathscr{L}$ over an open subset $V\subset X$ is the same thing as a section of $\mathscr{L}$ over $Z_{V}:=\pi^{-1}(V)$. The argument given (in the algebraic case) in [45, Tag 0D2M] shows that for $d_{0}=d_{0}(x)>0$ big enough, the evaluation map
(3.3) $E_{x}:\pi_{*}(\mathscr{L}^{\otimes d})_{x}\to H^{0}(Z_{x},\mathscr{L}_{x}^{\otimes d})$
is surjective for $d\geq d_{0}$. This means that every holomorphic section $\sigma_{x}:Z_{x}\to L_{x}^{\otimes d}$ of the line bundle $L_{x}^{\otimes d}\to Z_{x}$ extends to a holomorphic section $\sigma$ of $L^{\otimes d}|{Z_{V}}\to Z_{V}$ for an open neighbourhood $V$ of $x$. Denote by $\sigma]_{x}\in\pi_{*}(\mathscr{L}^{\otimes d})_{x}$ the associated element of the stalk $\pi_{*}(\mathscr{L}^{\otimes d})_{x}$, so $E_{x}([\sigma]_{x})=\sigma_{x}$. Since the sheaf $\pi_{*}\mathscr{L}^{\otimes d}$ is $\mathscr{O}_{X}$-coherent, Cartan’s Theorem A [27, p. 124] shows that $\pi_{*}(\mathscr{L}^{\otimes d})_{x}$ is generated as an $\mathscr{O}_{X,x}$-module by global sections of $\pi_{*}(\mathscr{L}^{\otimes d})$. Hence, there are sections $\xi_{1},\ldots,\xi_{m}\in\Gamma(X,\pi_{*}(\mathscr{L}^{\otimes d}))$ and germs of holomorphic functions $g_{1},\ldots,g_{m}\in\mathscr{O}_{X,x}$ such that
(3.4) $\sigma]_{x}=\sum_{j=1}^{m}g_{j}[\xi_{j}]_{x}\in\pi_{*}(\mathscr{L}^{\otimes d})_{x}.$
Since $X$ is Stein, there are functions $\tilde{g}_{j}\in\mathscr{O}(X)$ such that $\tilde{g}_{j}(x)=g_{j}(x)$ for $j=1,\ldots,m$. Then, $\tilde{\sigma}:=\sum_{j=1}^{m}\tilde{g}_{j}\xi_{j}\in\Gamma(X,\pi_{*}(\mathscr{L}^{\otimes d}))$ is a section whose value at $x$ equals
$E_{x}(\tilde{\sigma})=\sum_{j=1}^{m}\tilde{g}_{j}(x)E_{x}(\xi_{j})=\sum_{j=1}^{m}g_{j}(x)E_{x}(\xi_{j})=E_{x}([\sigma]_{x})=\sigma_{x}$
(see (3.4)). This completes the proof. ∎
Remark 3.2.
The integer $d_{0}=d_{0}(x)>0$ in the above argument may depend on $x$. For this reason, we do not know whether there is a proper embedding $F:\Omega\to X\times\mathbb{C}^{n}$ of class $\mathscr{A}(\Omega)$ and of the form (3.1) for some $n\in\mathbb{N}$. However, such an embedding exists and can be chosen fully algebraic if the base $X$ is affine algebraic due to finiteness of the above procedure, using the Zariski topology as opposed to the Hausdorff topology. The following is a special case of a lemma due to R. van Dobben de Bruyn [13]. This result will not be used in the paper and is stated only for the sake of comparison.
Lemma 3.3.
Assume that $X$ is an affine algebraic variety, $Z$ is an algebraic variety, $\pi:Z\to X$ is a proper algebraic submersion with one dimensional fibres, and $s_{1},\cdots,s_{m}$ for $m\geq 1$ are algebraic sections. If all fibres $Z_{x}\setminus\bigcup_{i=1}^{m}s_{i}(x)$ $(x\in X)$ are affine then there exists a proper algebraic morphism $F:\Omega=Z\setminus\bigcup_{i=1}^{m}s_{i}(X)\to X\times\mathbb{C}^{n}$ of the form (3.1) for some $n\in\mathbb{N}$.
Recall (see e.g. [33]) that a compact set $K$ in a complex manifold $\Omega$ is said to be $\mathscr{O}(\Omega)$-convex if it equals its $\mathscr{O}(\Omega)$-convex hull
$\widehat{K}_{\mathscr{O}(\Omega)}=\{z\in\Omega:|f(z)|\leq\sup_{K}|f|\ \text{for all}\ f\in\mathscr{O}(\Omega)\}.$
The following Oka–Weil type approximation theorem for the Stein domain $\Omega=Z\setminus H$ in Theorem 3.1 shows that the algebra $\mathscr{A}(\Omega)$ is dense in $\mathscr{O}(\Omega)$ in the compact-open topology.
Theorem 3.4.
Let $\pi:Z\to X$, $H\subset Z$, and $\Omega=Z\setminus H$ be as in Theorem 3.1. Given a compact $\mathscr{O}(X)$-convex set $K\subset\Omega$, every holomorphic function on a neighbourhood of $K$ is a uniform limit on $K$ of functions in $\mathscr{A}(\Omega)$. This holds in particular for the Teichmüller family $\Omega=V(g,n)$ (1.1) for any $g\geq 0$ and $n\geq 1$.
Proof.
Choose a relatively compact Runge domain $U\Subset X$ such that $\pi(K)\subset U$. Let $F:\Omega\to X\times\mathbb{C}^{n}$ be a map of class $\mathscr{A}(\Omega)$, furnished by Theorem 3.1 (see (3.1)), such that $F:\Omega_{U}\to U\times\mathbb{C}^{n}$ is a proper embedding. Its image $F(\Omega_{U})$ is then a closed complex submanifold of the Stein manifold $U\times\mathbb{C}^{n}$. It follows that the set $F(K)$ is holomorphically convex in $U\times\mathbb{C}^{n}$, and hence in $X\times\mathbb{C}^{n}$ since $U$ is Runge in $X$. Let $f$ be a holomorphic function on a Stein neighbourhood $V\subset\Omega$ of $K$. We may assume that $\pi(V)\subset U$. Then, $\tilde{f}=f\circ F^{-1}:F(V)\to\mathbb{C}$ is a holomorphic function on the open subset $F(V)$ in the closed complex submanifold $F(\Omega_{U})$ of $U\times\mathbb{C}^{n}$. By the Cartan–Oka extension theorem, $\tilde{f}$ extends to a holomorphic function on a neighbourhood of $F(K)$ in $U\times\mathbb{C}^{n}$. (This can also be seen by precomposing $\tilde{f}$ by a holomorphic retraction from a neighbourhood of $F(\Omega_{U})$ in $U\times\mathbb{C}^{n}$ onto $F(\Omega_{U})$, given by Docquier and Grauert [14]; see also [19, Theorem 3.3.3, p. 74].) By the Oka–Weil theorem, $\tilde{f}$ can be approximated uniformly on $F(K)$ by functions $g\in\mathscr{O}(X\times\mathbb{C}^{n})$. Every such function is of the form $g(x,w)=\sum_{\alpha\in\mathbb{Z}_{+}^{n}}g_{\alpha}(x)w^{\alpha},$ where $w=(w_{1},\ldots,w_{n})$ are coordinates on $\mathbb{C}^{n}$ and $g_{\alpha}\in\mathscr{O}(X)$ for every $\alpha\in\mathbb{Z}_{+}^{n}$. By cutting off the power series at a finite level gives functions $\mathscr{A}(X\times\mathbb{C}^{n})$ (polynomials in $w$ depending holomorphically on $x\in X$) approximating $g$ on $F(K)$. Precomposing such functions with $F$ gives the desired approximation of $f$ by functions in $\mathscr{A}(\Omega)$. ∎
The following is an immediate corollary to Theorem 3.4.
Corollary 3.5.
(Assumptions as in Theorem 3.4.) A compact set $K$ in $\Omega$ is $\mathscr{O}(\Omega)$-convex if and only if it is $\mathscr{A}(\Omega)$-convex.
It follows that the functions in $\mathscr{A}(\Omega)$ satisfy the axioms of a Stein manifold, i.e., they separate points in $\Omega$ and provide holomorphic convexity.
4. A relative Oka principle for fibrewise algebraic maps
In this section, we describe a relative Oka principle (see Theorem 4.1) for holomorphic fibrewise algebraic maps from the Teichmüller family $V(g,n)$ $(n\geq 1)$, or more generally from any manifold $\Omega=Z\setminus H$ as in Theorem 3.1, to a class of complex algebraic manifolds which includes all flexible manifolds in the sense of Arzhantsev et al. [9]. In order to state this result, we must recall the following notions; see Gromov [29, 0.5, p. 8.5.5], [19, Definition 5.6.13], or [20, Definition 3.1].
A dominating holomorphic spray on a complex manifold $Y$ is a holomorphic map $s:E\to Y$ from the total space of a holomorphic vector bundle $E\to Y$ such that
(4.1) $s(0_{y})=y\ \ \text{and}\ \ ds_{0_{y}}(E_{y})=T_{y}Y\ \ \text{for every $y\in Y$.}$
Here, $0_{y}$ is the origin of the fibre $E_{y}$ of $E$ over the point $y\in Y$, and $E_{y}$ is considered as a $\mathbb{C}$-linear subspace of the tangent space $T_{0_{y}}E$. (Identifying $Y$ with the zero section of $E$, we have $TE|Y\cong TY\oplus E$.) The spray is algebraic if the vector bundle $E\to Y$ and the map $s:E\to Y$ are algebraic. A complex manifold $Y$ is elliptic if it admits a dominating holomorphic spray, and is special elliptic if such a spray exists on a trivial bundle $E=Y\times\mathbb{C}^{m}$ for some $m\in\mathbb{N}$. An algebraic manifold $Y$ is algebraically (special) elliptic if these conditions hold with algebraic sprays.
An algebraic manifold $Y$ is said to be flexible [9] if its tangent bundle $TY$ is pointwise spanned by finitely many LND’s, that is, algebraic vector fields $V_{j}$ with complete algebraic flows $\phi_{t}^{j}$ $(t\in\mathbb{C},\ j=1,\ldots,m)$. The map $s:Y\times\mathbb{C}^{m}\to Y$ defined by
$s(y,t_{1},\ldots,t_{m})=\phi^{1}_{t_{1}}\circ\cdots\phi^{m}_{t_{m}}(y)$
is a dominating algebraic spray on $Y$, so a flexible manifold is algebraically special elliptic. References to examples of flexible manifolds can be found in [20, p. 394].
The following result is an analogue of the relative Oka principle for algebraic maps from affine algebraic varieties to algebraically elliptic manifolds; see [18, Theorem 3.1], [19, Theorem 6.15.1], and [20, Theorem 6.4].
Theorem 4.1.
Let $\pi:Z\to X$, $H\subset Z$, and $\Omega=Z\setminus H$ be as in Theorem 3.1, and let $Y$ be an algebraically special elliptic manifold. Given a map $f:Z\to Y$ of class $\mathscr{A}(Z,Y)$ (i.e., $f$ is holomorphic and fibrewise algebraic), a compact $\mathscr{O}(\Omega)$-convex set $K\subset\Omega$, and a homotopy of holomorphic maps $f_{t}:U\to Y$ $(t\in[0,1)$ on an open neighbourhood $U$ of $K$ with $f_{0}=f|_{U}$, there are holomorphic maps $F:\Omega\times\mathbb{C}\to Y$ such that $F(\cdotp,0)=f$ and $F(\cdot,t)\in\mathscr{A}(\Omega,Y)$ approximates $f_{t}$ as closely as desired uniformly on $K$ and uniformly in $t\in[0,1]$. In particular, a holomorphic map $\Omega\to Y$ that is homotopic to a map in $\mathscr{A}(\Omega,Y)$ is a limit of maps in $\mathscr{A}(\Omega,Y)$ uniformly on compacts. This holds in particular for the Teichmüller family $\Omega=V(g,n)$ for any $g\geq 0$ and $n\geq 1$.
Proof.
It suffice to inspect the proof of [18, Theorem 3.1] and apply Theorem 3.4. Let $p:E\to Y$ be an algebraic vector bundle and $s:E\to Y$ be a dominating algebraic spray. Consider the commuting diagram
[…]
where the map $\tilde{p}$ in the first column is the $f$-pullback of the vector bundle $p:E\to Y$, the map $\iota$ is the top row is the natural inclusion, and $\tilde{s}=s\circ\iota$ is a dominating holomorphic spray over $f$ (the pullback of $s$ by $f$). Note that $\tilde{s}=f$ holds on the zero section of $f^{*}E$. Since $f$ is fibrewise algebraic and $s$ is algebraic, the maps $\iota$ and $\tilde{s}$ are algebraic on $f^{*}E|\Omega_{x}$ for every $x\in X$.
Replacing the spray bundle $(E,p,s)$ with a sufficiently high iterate of itself (see [19, Definition 6.3.5] for this notion) and shrinking $U$ around $K$ if necessary, [19, Proposition 6.5.1] give a lift of the homotopy of holomorphic maps $f_{t}:U\to Y$ in the theorem to a homotopy of holomorphic sections $\xi_{t}:U\to f^{*}E|_{U}$, with $\xi_{0}$ the zero section, such that $\tilde{s}\circ\xi_{t}=f_{t}$ for $t\in[0,1]$.
Assume now that $p:E\to Y$ is a trivial algebraic vector bundle. Then, every iterate of $(E,p,s)$ is also a trivial bundle. Since $f:\Omega\to Y$ is fibrewise algebraic, the pullback bundle $f^{*}E\to\Omega$ is fibrewise algebraically trivial. Hence, sections of $f^{*}E\to\Omega$ can be identified with maps from $\Omega$ to the fibre, which is a Euclidean space. Thus, Theorem 3.4 lets us approximate every section $\xi_{t}$ uniformly on $K$ by a holomorphic fibrewise algebraic section $\tilde{\xi}_{t}:\Omega\to f^{*}E$. The map $F_{t}=\tilde{s}\circ\tilde{\xi}_{t}:\Omega\to Y$ is then of class $\mathscr{A}(\Omega,Y)$ and it approximates $f_{t}$ on $K$ for every $t\in[0,1]$. The approximation can be made uniform in $t\in[0,1]$ by including $[0,1]$ in a standard way into $\mathbb{C}$ and applying Theorem 3.4 with the compact holomorphically convex set $K\times[0,1]$ in $\Omega\times\mathbb{C}$. (The homotopy $\{\xi_{t}\}_{t\in[0,1]}$ must first be approximated on $K\times[0,1]$ by a holomorphic map in a neighbourhood of $K\times[0,1]$ in $\Omega\times\mathbb{C}$, which is an application of Mergelyan’s theorem.) This gives a holomorphic map $F:\Omega\times\mathbb{C}\to Y$ satisfying the conclusion of the theorem which is algebraic on $\Omega_{x}\times\mathbb{C}$ for every $x\in X$. ∎
Remark 4.2.
The relative algebraic Oka principle in [18, Theorem 3.1] holds for all algebraically elliptic target manifolds $Y$. (It was stated for the ostensibly bigger class of algebraically subelliptic manifolds, but it has recently been shown by Kaliman and Zaidenberg [36] that these two classes of manifolds coincide. See also [20, Theorem 6.2] for a more complete list of equivalent algebraic ellipticity conditions.) We do not know whether Theorem 4.1 holds for this bigger class of manifolds.
An example of a flexible manifold, which is of particular interest in the theory of minimal surfaces in Euclidean spaces $\mathbb{R}^{n}$, $n\geq 3$, is the punctured null quadric
$\mathbf{A}=\{z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}\setminus\{0\}:z_{1}^{2}+z_{2}^{2}+\cdots+z_{n}^{2}=0\}.$
(See [4, Proposition 1.15.3] and [8].) Algebraic 1-forms with values in $\mathbb{A}$ on finitely punctured compact Riemann surfaces, which satisfy the real period vanishing conditions on closed curves in $M$, gives rise to conformal minimal surfaces $M\to\mathbb{R}^{n}$ of finite total curvature via the Enneper–Weierstrass representation formula, and vice versa. (See [40] and [4, Chap. 4] for a discussion of this topic.) Hence, the fact that the Teichmüller family $V(g,n)$ for $n\geq 1$ admits nontrivial fibrewise algebraic maps $V(g,n)\to\mathbf{A}$ raises the question whether there one can realise the fibres of $V(g,n)$ as a family of minimal surfaces of finite total curvature depending holomorphically on the parameter $t\in T(g,n)$.
Acknowledgements. Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia. I wish to thank Finnur Lárusson for asking the question answered by Theorem 1.1 and for proposing Corollary 1.3 in a private communication, and Daniel Huybrechts and Vladimir Marković for useful remarks and suggestions.